mark is y years old. in five years time aoife will be twice as old as mark and daniel will be two years younger than aoife is now. write an expression in y for the sum of the 3 ages now. show workings please :)
Answer by ankor@dixie-net.com(22740)

mark is y years old.
In five years time Aoife will be twice as old as Mark
A + 5 = 2(y+5)
A + 5 = 2y + 10
A = 2y + 10 - 5
A = 2y + 5
:
and Daniel will be two years younger than Aoife is now.
d + 5 = A - 2
d = A - 2 - 5
d = A - 7
Replace A with (2y+5)
d = (2y+5) - 7
d = 2y - 2
:
write an expression in y for the sum of the 3 ages now.
Sum = y + A + d
replace A and d with expressions derived above:
S = y + (2y+5) + (2y-2)
S = y + 4y + 3
S = 5y + 3 is the sum of the three ages